OpenStudy (anonymous):

1. Suppose that the amount of algae in a pond doubles every 4 hours. If the pond initially contains 90 pounds of algae, how much algae will be in the pond after 12 hours? (1 point) would it be 720lbs

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

12 hours is 3 doubling periods (since they are 4 hours each) so your 90 pounds doubles 3 times

OpenStudy (anonymous):

in four hours you have \(90\times 2=180\) in another 4 hours you have \(180\times 2=360\) and in another 4 hours you have \(360\times 2=720\) so yes, you are right

OpenStudy (anonymous):

so times 90 3times? i get 729000

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

2. Suppose that the diameters of different types of hairs are 1.7x 10^-4cm, 1.4 x 10^-3cm, 1.2 x 10^-5cm, and 1.9 x 10^-4cm? What is the order of the hairs from least to greatest diameters? (1 point) CHECK MY ANSWER PLEASE 1.2 x 10^-5, 1.7 x 10^-4, 1.9 x 10^-4, 1.4 x 10^-3